# Weichuang Shao Iman Yi Liao Tomas Henrique Bode Maul Tissa Chandesa

### v1

@@ 4.4 Computational Efficiency
By removing the dependence on pre-training, AMAuT achieves remarkable computational efficiency. The training of AMAuT can be completed entirely on a single NVIDIA RTX 4090 GPU, while maintaining performance comparable to or better than several leading pre-trained models such as SSAST [16], EAT [6], HuBERT [19, 27], and Qwen-Audio [8] (see Table 9).
Table 9: Comparison of pre-training and fine-tuning hours for AMAuT and state-of-the-art pre-trained models across multiple datasets. Pre-training and fine-tuning durations are reported for each algorithm, where available, from their respective original publications.
@@
SSAST [16] SC1 10 days [16] unreported
@@
EAT [6] SC2 58h [6] unreported
AST [14] SC2 3 days [35] unreported
PDC [7] SC2 unreported unreported
@@
HuBERT [19, 27] AM 137.75 h [19] unreported
XLS-R [2, 27] AM unreported unreported
Wav2Vec2 [3, 27] AM 7.5 days [3] unreported
@@
Qwen-Audio [8] VS unreported unreported
EfficientNet [17, 34] VS unknown unreported
@@
Audio Flamingo [25] CS unreported unreported
Qwen-Audio [8] CS unreported unreported
As shown in Table 9, SSAST [19] requires four NVIDIA GTX Titan X GPUs, 1.95 million AudioSet [13] samples, and 281 thousand LibriSpeech [30] samples for pre-training, a process that takes approximately 10 days. EAT [6] performs pre-training on four RTX 3090 GPUs using the AudioSet [13], requiring about 58 hours. AST [14] does not perform self-pre-training but instead reuses weights from DeiT [35], which itself was trained on ImageNet [32] using four 16 GB V100 GPUs for three days. HuBERT [19, 27] needs 32 GPUs and a total of 60,960 hours of audio (960 hours from LibriSpeech [30] plus 60,000 hours from Libri-Light [22]) for pre-training, which takes approximately 137.75 hours. Similarly, XLS-R [2, 27] requires 128 GPUs and several large-scale multilingual datasets, including VoxPopuli [40], Multilingual LibriSpeech [31], CommonVoice [1], VoxLingua107 [37], and BABEL [12]. Wav2Vec 2.0 [3, 27] also needs 128 V100 GPUs, running for 2.3 days on LibriSpeech [30] and 5.2 days on Libri-Light [22].
Although the pre-training durations and hardware requirements of Qwen-Audio [8] are not explicitly reported, its large-scale use of 30 audio datasets (covering speech, sound, and music) suggests that its pre-training would also be highly resource-intensive. Likewise, Audio Flamingo [25] indicates the use of eight NVIDIA A100 GPUs for pre-training but does not disclose the total compute time or energy cost.
Table 10: Comparison of GPU hours for AMAuT and state-of-the-art pre-trained models across multiple datasets. GPU hours = number of GPUs $\times$ (pre-training + training time). Pre-training and fine-tuning durations are reported for each algorithm where available. Data are collected from original publications unless otherwise noted.
@@ 6 Limitations and Future Work
    Performance is sensitive to the Mel-spectrogram [36, 20] configuration and TTDA loss parameters (Eq. 5). The absence of universal tuning guidelines indicates a need for automated hyperparameter search in future work.
@@ 7 Conclusion
Consequently, AMAuT can be trained from scratch on diverse datasets. Evaluations on five benchmark datasets, such as AM, SC1, SC2, VS, and CS, demonstrate that AMAuT achieves performance comparable to state-of-the-art pre-trained models like SSAST [16], EAT [6], HuBERT [19], Qwen-Audio [8], and Audio Flamingo [25], with test accuracies of 99.82%, 96.31%, 96.21%, 92.26%, and 72.58%, respectively. Notably, this performance is achieved using under 3% of the GPU time required by the compared models.
@@ A.2 Test-time Domain Adaptation (TTDA)
  • Nuclear-Norm Maximization (NM) [26]:
@@
  • Entropy Minimization (EN) [41]:
@@
  • A modified Generalized Entropy (GEN) [23]:
@@ A.4 Training Loss
To prevent overconfidence and improve generalization, this study leverages cross-entropy and label-smoothing regularization (LSR) [33] for training loss:
@@
Appendix B Hyperparameter Settings
B.1 TTDA Hypermeters
Table 11: Test-time domain adaptation hyperparameter settings
Set $\alpha$ $\beta$ $\gamma$ q
AM 1.0 0.5 0.5 1.1
SC1 0.2 1.0 0.0 0.8
SC2 0.0 1.0 0.2 1.1
VS 1.0 0.5 0.5 1.1
CS 1.0 0.5 0.5 1.1
B.2 Learning Rate Scheduler Hyperparameters
Table 12: Learning rate scheduler hyperparameter settings
Set lr $\lambda$ $\eta$
AM $10^{-3}$ 10 40
SC1 $10^{-3}$ 10 50
SC2 $10^{-3}$ 10 50
VS $10^{-3}$ 10 80
CS $10^{-3}$ 10 40

### v2

@@ 4.4 Computational Efficiency
By removing the dependence on pre-training, AMAuT achieves remarkable computational efficiency. The training of AMAuT can be completed entirely on a single NVIDIA RTX 4090 GPU, while maintaining performance comparable to or better than several leading pre-trained models such as SSAST [gong2022ssast], EAT [chen2024eat], HuBERT [hsu2021hubert, la2024benchmarking], and Qwen-Audio [chu2023qwen] (see Table 10).
Table 10: Comparison of pre-training and fine-tuning hours for AMAuT and state-of-the-art pre-trained models across multiple datasets.
@@
SSAST [gong2022ssast] SC1 10 days [gong2022ssast] unreported
@@
EAT [chen2024eat] SC2 58h [chen2024eat] unreported
AST [gong2021ast] SC2 3 days [touvron2021training] unreported
PDC [chrysos2022augmenting] SC2 unreported unreported
@@
HuBERT [hsu2021hubert, la2024benchmarking] AM 137.75 h [hsu2021hubert] unreported
XLS-R [babu2021xls, la2024benchmarking] AM unreported unreported
Wav2Vec2 [baevski2020wav2vec, la2024benchmarking] AM 7.5 days [baevski2020wav2vec] unreported
@@
Qwen-Audio [chu2023qwen] VS unreported unreported
EfficientNet [gong_vocalsound, tan2019efficientnet] VS unknown unreported
@@
Audio Flamingo [kong2024audio] CS unreported unreported
Qwen-Audio [chu2023qwen] CS unreported unreported
  • 1
    Pre-training and fine-tuning durations are reported for each algorithm, where available, from their respective original publications.
As shown in Table 10, SSAST [gong2022ssast] requires four NVIDIA GTX Titan X GPUs, 1.95 million AudioSet [gemmeke2017audio] samples, and 281 thousand LibriSpeech [panayotov2015librispeech] samples for pre-training, a process that takes approximately 10 days. EAT [chen2024eat] performs pre-training on four RTX 3090 GPUs using the AudioSet [gemmeke2017audio], requiring about 58 hours. AST [gong2021ast] does not perform self-pre-training but instead reuses weights from DeiT [touvron2021training], which itself was trained on ImageNet [ILSVRC15] using four 16 GB V100 GPUs for three days. HuBERT [hsu2021hubert, la2024benchmarking] needs 32 GPUs and a total of 60,960 hours of audio (960 hours from LibriSpeech [panayotov2015librispeech] plus 60,000 hours from Libri-Light [kahn2020libri]) for pre-training, which takes approximately 137.75 hours. Similarly, XLS-R [babu2021xls, la2024benchmarking] requires 128 GPUs and several large-scale multilingual datasets, including VoxPopuli [wang2021voxpopuli], Multilingual LibriSpeech [pratap2020mls], CommonVoice [ardila2019common], VoxLingua107 [valk2021voxlingua107], and BABEL [gales2014speech]. Wav2Vec 2.0 [baevski2020wav2vec, la2024benchmarking] also needs 128 V100 GPUs, running for 2.3 days on LibriSpeech [panayotov2015librispeech] and 5.2 days on Libri-Light [kahn2020libri].
Although the pre-training durations and hardware requirements of Qwen-Audio [chu2023qwen] are not explicitly reported, its large-scale use of 30 audio datasets (covering speech, sound, and music) suggests that its pre-training would also be highly resource-intensive. Likewise, Audio Flamingo [kong2024audio] indicates the use of eight NVIDIA A100 GPUs for pre-training but does not disclose the total compute time or energy cost.
Table 11: Comparison of GPU hours for AMAuT and state-of-the-art pre-trained models across multiple datasets.
@@
  • 1
    GPU hours = number of GPUs $\times$ (pre-training + training time).
  • 2
    Pre-training and fine-tuning durations are reported for each algorithm where available.
  • 3
    Data are collected from original publications unless otherwise noted.
@@ 6 Limitations and Future Work
    Performance is sensitive to the Mel-spectrogram [ustubioglu2023mel, hwang2020mel] configuration and TTDA loss parameters (Eq. 5). The absence of universal tuning guidelines indicates a need for automated hyperparameter search in future work.
@@ 7 Conclusion
Consequently, AMAuT can be trained from scratch on diverse datasets. Evaluations on five benchmark datasets, such as AM, SC1, SC2, VS, and CS, demonstrate that AMAuT achieves performance comparable to state-of-the-art pre-trained models like SSAST [gong2022ssast], EAT [chen2024eat], HuBERT [hsu2021hubert], Qwen-Audio [chu2023qwen], and Audio Flamingo [kong2024audio], with test accuracies of 99.82%, 96.31%, 96.21%, 92.26%, and 72.58%, respectively. Notably, this performance is achieved using under 3% of the GPU time required by the compared models.
@@ A.2 Test-time Domain Adaptation (TTDA)
  • Nuclear-Norm Maximization (NM) [kumar2023conmix]:
@@
  • Entropy Minimization (EN) [wang2021tent]:
@@
  • A modified Generalized Entropy (GEN) [kim2023sgem]:
@@ A.4 Training Loss
To prevent overconfidence and improve generalization, this study leverages cross-entropy and label-smoothing regularization (LSR) [szegedy2016rethinking] for training loss:
@@
Appendix B Consistency Analysis
To evaluate the consistency of independently trained AMAuT models, an agreement rate is computed based on the alignment of their predictions. Let the confusion matrix between two prediction sets $p_{1}$ and $p_{2}$ be defined as:
$cm=\text{confusion\_matrix}(p_{1},p_{2})$ (14)
The overall agreement rate is then calculated as:
$\text{agreement rate}=\frac{\sum_{i}cm_{i,i}}{\sum_{i,j}cm_{i,j}}$ (15)
where $cm_{i,i}$ denotes the number of samples for which both predict the same class label $i$, and $\sum_{i,j}cm_{i,j}$ represents the total number of samples considered.
A higher agreement rate indicates greater consistency between models trained independently under the same configuration, suggesting stable convergence and robust generalization of the AMAuT framework.
Appendix C Hyperparameter Settings
Table 12: Hyperparameter Settings
Set $\alpha$ $\beta$ $\gamma$ $q$ lrlr $\lambda$ $\eta$
AM 1.0 0.5 0.5 1.1 $10^{-3}$ 10 40
SC1 0.2 1.0 0.0 0.8 $10^{-3}$ 10 50
SC2 0.0 1.0 0.2 1.1 $10^{-3}$ 10 50
VS 1.0 0.5 0.5 1.1 $10^{-3}$ 10 80
CS 1.0 0.5 0.5 1.1 $10^{-3}$ 10 40
  • 1
    The parameters, $\alpha$, $\beta$, and $\gamma$, correspond to the weighting coefficients in Eq. 5.
  • 2
    The coefficient $q$ denotes the scaling factor in Eq. 9.
  • 3
    The learning rate lrlr, adjustment factor $\lambda$, and scheduling cardinality $\eta$ follow the update rule defined in Eq. 6.